Find the irreducible polynomial of 1 + i over Q.
Answer #1

\large \\ \text{ Clearly , } 1+i\notin \mathbb Q\\ \text{ So there is no linear irreducible polynomial exists. The irreducible polynomial must have atleast degree 2}\\ x=1+i\\ (x-1)^2=-1\implies (x-1)^2+1=0\\ \text{ So } (x-1)^2+1=x^2-2x+2=0\\ \text{ So root of above equation is } 1+i \\ \text{ and we had find 2nd order polynomial which is force to be irreducible. Even rational root test support this .}\\\large \\ \text{ Rational Root test: Given}\\ a_nx^n+a_{n-1}x^{n-1}+....+a_0=0\\ \text{ with integer } a_i ;a_n\neq 0 \text{ Then every rational root of given polynomial is of forms } \frac{\pm \text{Factors of }a_0}{\pm \text{Factors of }a_n}\\\\ a_0=2,a_2=1\\\\ \text{ Divisor of } a_0=\pm1,\pm2\\\\ \text{Divisor of } a_2=\pm1\\\\ \text{We will check following rational numbers } \pm \frac{1,2}{1}\\\\ \text{ All possible roots are } : 1,-1,2,-2\\\\ p(x)=x^2-2x+2\\ p(1)=1,p(-1)=5,p(2)=2,p(-2)=10\\ \text{ So polynomial p(x) is irreducuble over } \mathbb Q.\\\\ \textbf { So irreducible polynomial of } 1+i \textbf{ is } x^2-2x+2 \textbf{ over } \mathbb Q